Plants consume carbon dioxide gas as they produce sugar during photosynhesis


                       6CO2 +6H2O → C6H12O6 + 6O2

To produce 50g of sugar, C6H12O6  how many litres of  gas at SATP does a sugar beet need to consume? [T/I-7]




1
Expert's answer
2022-04-12T15:38:03-0400

Moles of sugar = 50 / 180 = 0.278 (mol)

Six more moles of gas (CO2) is consumed: moles CO2 = 6 * 0.278 = 1.67 (mol)

Volume = 1.67 * 22.4 = 37.3 (L)